OpenStudy (anonymous):

[How would I check a formula for dimensional consistency? e.g., T= 2π √L/G is the formula for a single pendulum to complete a single oscillation. T = Time to complete oscillation L = Length of the pendulum G = Acceleration due to gravity

OpenStudy (aum):

2 and π are dimensionless and so we don't have to worry about it. \[ \sqrt{\frac LG } = \sqrt{\frac {m}{m/s^2}} = \sqrt{s^2} = s \]

OpenStudy (aum):

Right hand side reduces to seconds (or time) and the left hand side is Time period in seconds. So they formula is dimensionally consistent.

OpenStudy (anonymous):

Wow. You actually made this make sense to me. I completely understand it now. Thanks so much!

OpenStudy (aum):

You are welcome. Glad to be able to help.